OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5522 — boi-1-3k
Incident Report for Web Cloud
Resolved
We have detected an optical problem:

TE0/2: Temperature high alarm; Operating value: 113.9 C, Threshold value: 100.0 C.

A technician is underway to replace the equipment. An interuption of a few minutes is expected during maintenance.



Update(s):

Date: 2011-06-14 17:34:31 UTC
The second optical bloc is replaced. We renforce the monitoring
of this equipment for the next days in order to be sure of its
stability. The working threshold are currently normal.

Date: 2011-06-14 17:32:55 UTC
We are replacing the optical bloc.

Date: 2011-06-14 17:32:30 UTC
The first optical bloc faces itself problems.

A thechnician is on the road to replace it.
Posted Jun 14, 2011 - 11:49 UTC